# HyperSIGMA: Hyperspectral Intelligence Comprehension Foundation Model

### v1

@@ 5.7.1 Model Scalability
Loading Different Weights In HyperSIGMA, both the spatial and spectral subnetworks utilize a ViT structure, allowing pre-trained weights from one branch to be reused in the other. We explored this by fine-tuning on the HanChuan dataset, initializing SpatSIGMA and HyperSIGMA with different pre-training weights, as detailed in Table XVI. Using spectral ViT pre-training weights in SpatSIGMA led to a performance drop, highlighting the necessity of spatial pre-training (i.e., matched) weights for the spatial subnetwork. HyperSIGMA also showed reduced accuracy with mismatched weights, with the worst performance occurring when both subnetworks used mismatched weights. Comparing SpatSIGMA-Spec and HyperSIGMA-DualSpec, we found that matched weights in the spectral branch did not enhance accuracy if the spatial branch weights were mismatched. Conversely, comparing SpatSIGMA and HyperSIGMA-DualSpat revealed that mismatched spectral branch weights caused a milder performance decline. These results indicate that the spatial branch has stricter weight initialization requirements. Ultimately, SpatSIGMA and HyperSIGMA performed best with matched weights, underscoring the importance of structure-weight matching and confirming the auxiliary role of the spectral subnetworks.
5.7.2 Model Robustness
Figure 23: Overall accuracies of different methods with limited training samples: (a) HanChuan dataset, (b) HongHu dataset.
TABLE XVII: OAs (%) of different methods under various attacks on the Indian Pines dataset. Best and 2nd-best results are highlighted.
Method SSFCN SACNet RCCA S3ANet SpatSIGMA HyperSIGMA
Clean Test Set 82.57 88.63 94.51 96.23 96.28 96.76
FGSM 60.26 80.19 91.33 95.48 96.19 96.62
PGD 1.48 30.18 82.96 42.01 95.23 95.92
Table XVII presents the classification accuracy on the original clean testing set and under various attacks. Our models show minimal accuracy decreases and consistently outperform others, regardless of the attack type. Compared to SpatSIGMA, HyperSIGMA, which leverages spectral information, achieves even higher accuracy. To further assess model stability, we increase the FGSM perturbation budget $\epsilon$ and plot the accuracy changes in Figure 24. Most comparison methods’ accuracies decline rapidly when $\epsilon$ reaches 0.4. In contrast, our models maintain stability, with only slight accuracy changes even at $\epsilon$ of 1.0. These results clearly demonstrate the robustness of our models against adversarial attacks.
5.7.3 Cross-modal Transferability
Figure 26: Visualization of detected oil leakage in various Gulf of Mexico regions by our models: (a) GM07, (b) GM13, (c) GM17, and (d) GM18.
Given the similarity between HSIs and multispectral images, we hypothesized that HyperSIGMA could be effectively applied to multispectral tasks. To test this, we conducted experiments on the OSCD dataset [244], which uses Sentinel-2 images for multispectral change detection (CD). The structure of HyperSIGMA for multispectral CD is similar to that shown in Fig. 15, but incorporates a simple feature pyramid network (FPN) like [245] to upsample features layer by layer, gradually recovering spatial details for the final prediction. We opted for a simple approach to highlight HyperSIGMA’s feature extraction capabilities, avoiding the influence of a complex change decoder. Data processing followed the protocol in [3]. Table XVIII compares HyperSIGMA’s performance with leading CD methods, including FC-EF [246], FC-Siam-Diff [246], FC-Siam-Conc [246], SiamCRNN [247], SNUNet [248], DSIFN [249], BIT [199], ChangeFormer [250], and ChangeMamba [251]. We also compared our method with SpectralGPT [3], a recent model tailored for multispectral data. HyperSIGMA surpassed the state-of-the-art ChangeMamba, achieving an F1 score of 58.53% versus 57.20%. Further enhancement using spectral information boosted HyperSIGMA’s performance to an F1 score of 59.28%, marking a 4.99% improvement over SpectralGPT.
TABLE XVIII: Change detection results of different methods on the OSCD dataset. SpectralGPT accuracy values are sourced from [3]. Best and 2nd-best results are highlighted.
Method OA Kappa F1 Precision Recall
FC-EF [246] 94.80 45.10 47.83 49.67 46.12
FC-Siam-Diff [246] 94.06 45.26 48.38 43.91 53.85
FC-Siam-Conc [246] 94.55 47.57 50.43 47.60 53.62
SiamCRNN [247] 95.53 52.08 54.42 57.58 51.60
SNUNet [248] 93.68 39.70 43.03 40.29 46.16
DSIFN [249] 96.04 52.22 54.21 67.32 45.38
BIT [199] 94.88 43.82 41.19 50.56 38.66
ChangeFormer [250] 95.40 52.15 54.57 55.66 53.51
ChangeMamba [251]. 96.04 54.82 57.20 56.08 58.36
SpectralGPT[3] - - 54.29 52.39 57.20
SpatSIGMA 96.08 56.49 58.53 64.59 53.50
HyperSIGMA 95.78 57.06 59.28 59.12 59.45
5.7.4 Real-world Applicability
6 Conclusion
In this paper, we present HyperSIGMA, the first hyperspectral foundation model with over 1 billion parameters for hyperspectral image (HSI) interpretation. Furthermore, we curate HyperGlobal-450K, the largest hyperspectral dataset to date, composed of HSIs from around the world, establishing a solid basis for self-supervised pre-training research. To address HSI redundancy, we propose a novel sparse sampling attention mechanism, enabling adaptive perception of relevant contextual regions with few learnable sampling points. We also design a spectral enhancement module to achieve effective spatial-spectral feature fusion. Comprehensive evaluations on both high-level and low-level hyperspectral tasks demonstrate HyperSIGMA’s superior performance. Further analysis reveals its excellent scalability, robustness, and cross-modal transferability, making it suitable for various real-world applications.
[…]

### v2

@@ I.2 More Results
Loading Different Weights In HyperSIGMA, both the spatial and spectral subnetworks utilize a ViT structure, allowing pre-trained weights from one branch to be reused in the other. We explored this by fine-tuning on the HanChuan dataset, and initializing SpatSIGMA and HyperSIGMA with different pre-training weights, as detailed in Table XXV. Using spectral ViT pre-training weights in SpatSIGMA led to a performance drop, highlighting the necessity of spatial pre-training (i.e., matched) weights for the spatial subnetwork. HyperSIGMA also showed reduced accuracy with mismatched weights, with the worst performance occurring when both subnetworks used mismatched weights. Comparing SpatSIGMA-Spec and HyperSIGMA-DualSpec, we found that matched weights in the spectral branch did not enhance accuracy if the spatial branch weights were mismatched. Conversely, comparing SpatSIGMA and HyperSIGMA-DualSpat revealed that mismatched spectral branch weights caused a milder performance decline. These results indicate that the spatial branch has stricter weight initialization requirements. Ultimately, SpatSIGMA and HyperSIGMA performed best with matched weights, underscoring the importance of structure-weight matching and confirming the auxiliary role of the spectral subnetworks.
Appendix K Cross-modal Transferability Experiments
Network Structure For collaborative classification, we adopt the same segmentation structure for HyperSIGMA following previous classification fine-tuning experiments, as shown in Fig. 14, where we input 128 $\times$ 128 patches, with the spatial patch size is set to 2. For the multispectral CD task, the structure of HyperSIGMA is similar to that shown in Fig. 26, but incorporates a simple feature pyramid network (FPN) like [149] to upsample features layer by layer, gradually recovering spatial details for the final prediction.
Experimental Details Data processing and label split of collaborative classification and CD tasks follow the protocol in [147] and [3], respectively.
Appendix L More Results of Oil Leakage Detection
Fig. 37 shows the visualization results of oil leakage detection using HyperSIGMA on other regions of the Gulf of Mexico.
Appendix M More Complexity Analyses and Experimental Details
Detailed Complexity Calculation Process Given a 2-D feature map $\mathbf{X}\in\mathbb{R}^{N\times D^{\prime}}$ that has $N$ tokens with dimensions in $D^{\prime}$, we first generate query, key and value maps $\mathbf{Q}=\{q_{1},\cdots,q_{N}\}$, $\mathbf{K}=\{k_{1},\cdots,k_{N}\}$, and $\mathbf{V}=\{v_{1},\cdots,v_{N}\}$, requiring a complexity of $O(3ND^{\prime 2})$. Then, we predict $N_{p}$ positions (including X and Y coordinates) for each query, spending $O(2NN_{p}D^{\prime})$ computational complexity. The new $D^{\prime}$ dimension key and value vectors in $\mathbf{K}^{\prime},\mathbf{V}^{\prime}\in\mathbb{R}^{N\times N_{p}\times D^{\prime}}$ are both sampled by the bilinear interpolation considering four nearest points, and this process totally costs $O(8NN_{p}D^{\prime})$ complexity. The multiplication $\otimes$ between $\mathbf{Q}$ and $\mathbf{K}^{\prime}$, and $\mathbf{Q}\otimes\mathbf{K}^{\prime}$ and $\mathbf{V}^{\prime}$, costing totally $O(2NN_{p}D^{\prime})$ computational complexity, while the softmax process in $\mathbf{Q}\otimes\mathbf{K}^{\prime}$ need a complexity of $O(NN_{p})$. Therefore, the computational complexity of SSA is $O(3ND^{\prime 2}+12NN_{p}D^{\prime}+NN_{p})$.
Experimental Settings We use the Indian Pines dataset, which contains $N=21,025$ pixels, and perform category prediction for each pixel via patch-level classification. For model input, we select a surrounding patch of size $p\times p$, resulting in an input size of $N\times p\times p$. We conduct experiments with different patch sizes $p$, maintaining a constant batch size of 8. To assess inference time, we measure the total time required to predict all pixels and then calculate the average time per pixel by dividing the total time by $N$.
Appendix N Datasheet for HyperGlobal-450K
N.1 Motivation
1. For what purpose was the dataset created?
A1: The HyperGlobal-450K dataset is designed to facilitate the development of large foundational models for HSI processing and analysis using self-supervised learning. Given the scarcity of large-scale, high-quality hyperspectral datasets with global coverage, this dataset integrates data from multiple sensors to build a model with powerful generalization ability. HyperGlobal-450K provides a comprehensive, diverse, and temporally extensive HSI resource, driving innovation in RS technologies and models, and supporting applications such as land cover classification and environmental monitoring.
2. Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
A2: The dataset was created by:
Di Wang (Wuhan University),
Meiqi Hu (Wuhan University),
Yao Jin (Wuhan University),
Yuchun Miao (Wuhan University),
Jiaqi Yang (Wuhan University),
Yichu Xu (Wuhan University),
Xiaolei Qin (Wuhan University),
Jiaqi Ma (Wuhan University),
Lingyu Sun (Wuhan University),
Chenxing Li (Wuhan University),
Chuan Fu (Chongqing University),
Hongruixuan Chen (University of Tokyo),
Chengxi Han (Wuhan University),
Naoto Yokoya (University of Tokyo),
Jing Zhang (Wuhan University),
Minqiang Xu (National Engineering Research Center of Speech and Language Information Processing),
Lin Liu (National Engineering Research Center for Speech and Language Information Processing),
Lefei Zhang (Wuhan University),
Chen Wu (Wuhan University),
Bo Du (Wuhan University),
Dacheng Tao (Nanyang Technological University),
Liangpei Zhang (Wuhan University).
3. Who funded the creation of the dataset?
A3: The dataset creation was funded by the affiliations of the authors involved in this work.
N.2 Composition
1. What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances(e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
A1: This dataset is composed of the image subsets captured by the EO-1 satellite and the Gaofen-5 satellite. In these subsets, each sample is a single HSI.
2. How many instances are there in total (of each type, if appropriate)?
A2: HyperGlobal-450K contains 447,072 HSIs, including 247,072 EO-1 image patches and 200,000 GF-5 samples.
3. Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
A3: HyperGlobal-450K is sampled from larger sets, which contains so far all images photoed by EO-1 and GF-5 satellites. For the EO-1 instance, since the locations are selected to cover the globe path/row pairs in the Worldwide Reference System as far as possible, it can represent the large EO-1 set to some extent. The GF-5 instances in HyperGlobal-450K can represent most of the large GF-5 set, since they encompass a variety of landscapes in China, such as Grasslands, forests, farmlands, etc.
4. What data does each instance consist of? “Raw” data (e.g., unprocessed text or images)or features? In either case, please provide a description.
A4: Each instance consists of a clipped HSI patch.
5. Is there a label or target associated with each instance? If so, please provide a description.
A5: No, the dataset is intended for self-supervised learning. Therefore, each instance is an individual RS image and does not possess any form of annotations.
6. Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
A6: No.
7. Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are drawn.
A7: Yes, the relationships between different instances are shown in filenames.
8. Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
A8: Yes. The entire dataset is designated for self-supervised learning algorithms. Therefore, we recommend the whole dataset be utilized for self-supervised learning research.
9. Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
A9: No.
10. Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a future user? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
A10: The dataset is self-contained as it is derived from raw satellite imagery downloaded from official sources, but users should refer to the original platforms for licensing terms and note that the availability of the raw data may vary over time.
11. Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals non-public communications)? If so, please provide a description.
A11: No.
12. Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
A12: No.
N.3 Collection process
1. How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
A1: The data associated with each instance is directly observable, as they are HSIs and can be acquired from EO-1 or GF-5 hyperspectral satellite imagery.
2. What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)? How were these mechanisms or procedures validated?
A2: The dataset consists of HSIs that were downloaded from USGS and CPEOS. The images were filtered based on cloud cover, location, and spectral bands. The cloud cover selection is performed online. The Python scripts are compiled to perform location and band selection. The satellite sensors were pre-calibrated, and all the scripts were thoroughly tested to ensure reliability.
3. If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
A3: The EO-1 instances were sampled from the aforementioned large set using three strategies: (1) Deterministic cloud cover selection: screening the data with cloud cover percentage less than 5% from 2011-2017 and excluding the data with unknown cloud cover percentage. (2) Random location selection: randomly selecting one image each year from each path/row location derived from the Worldwide Reference System (WRS). (3) Deterministic band selection: removing bad bands and water vapor absorption bands. For GF-5, the instances were sampled from related large sets by deterministic band and location selections.
4. Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)
A4: The data was collected and verified by the authors.
5. Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
A5: The data download takes approximately one month. Then, performing location selection costs about 1 day, while it takes about 7 days for band selection and clipping.
N.4 Preprocessing/cleaning/labeling
1. Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing 5 of missing values)? If so, please provide a description. If not, you may skip the remainder of the questions in this section.
A1: The EO-1 and GF-5B hyperspectral images are clipped into hyperspectral patches with a spatial size of 64×64, to facilitate network reading and training.
2. Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
A2: No.
3. Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
A3: We use Python for clipping.
N.5 Uses
1. Has the dataset been used for any tasks already? If so, please provide a description.
A1: No.
2. Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
A2: N/A.
3. What (other) tasks could the dataset be used for?
A3: HyperGlobal-450K can be used for unsupervised pre-training. Besides, it can be used for evaluation on specific tasks, such as image denoising and super-resolution. With further manual labeling, it can also be used for semantic segmentation, anomaly detection, target detection, etc.
4. Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a future user might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other undesirable harms (e.g., financial harms, legal risks) If so, please provide a description. Is there anything a future user could do to mitigate these undesirable harms?
A4: No.
5. Are there tasks for which the dataset should not be used? If so, please provide a description.
A5: No.
N.6 Distribution
1. Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
A1: Yes. The dataset will be publicly available.
2. How will the dataset be distributed (e.g., tarball on website, API, GitHub)?
A2: It will be publicly available on the project website at HyperGlobal-450K.
3. When will the dataset be distributed?
A3: The dataset will be distributed once the paper is accepted after peer review.
4. Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
A4: It will be distributed under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 License.
5. Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
A5: No.
6. Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
A6: No.
N.7 Maintenance
1. Who will be supporting/hosting/maintaining the dataset?
A1: The authors.
2. How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
A2: They can be contacted via email available on the project website.
3. Is there an erratum? If so, please provide a link or other access point.
A3: No.
4. Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
A4: No.
5. Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
A5: N/A.
6. If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to other users? If so, please provide a description.
A6: N/A.
Appendix O Per-class Accuracies on Hyperspectral Image Classification Datasets
We present the per-class classification accuracies of various methods on the Indian Pines, Pavia University, HanChuan, HongHu, Houston and ZY1-02D datasets in Tables XXVI-XXXI.
TABLE XXVI: Per-class classification accuracies of various methods on the Indian Pines dataset.
Class1 MSDN [60] SSFCN [61] FullyContNet [59] SpectralFormer [20] HSIC-FM [62] SSGRN [63] CSIL [64] IDCN [65] CLOLN [66] SpatSIGMA HyperSIGMA
1 55.08 88.89 82.41 84.26 92.59 100.00 99.07 86.02 81.46 73.80 44.28
2 52.60 19.56 50.73 36.84 34.39 53.87 42.38 57.50 67.76 87.62 91.99
3 38.72 29.96 57.76 39.35 27.28 32.07 62.03 53.49 49.78 81.31 88.96
4 37.13 54.04 75.92 51.54 45.23 86.78 87.22 81.67 52.68 73.19 72.18
5 61.44 43.62 78.65 63.57 61.66 77.80 80.20 84.25 95.84 96.69 100.00
6 75.77 51.02 73.52 64.31 41.76 63.19 88.75 86.94 98.35 94.00 88.25
7 21.11 83.33 92.59 100.00 55.56 100.00 100.00 97.43 45.07 65.00 100.00
8 86.78 87.75 92.52 82.19 66.17 100.00 99.72 88.91 97.78 98.28 92.88
9 11.77 96.67 100.00 96.67 100.00 100.00 100.00 100.00 46.72 15.62 8.06
10 38.53 51.42 78.52 50.49 16.53 69.75 64.00 71.36 84.64 74.00 89.56
11 61.94 33.44 64.16 42.52 21.06 69.12 55.01 60.35 75.00 87.04 87.23
12 39.16 30.30 57.06 29.10 34.08 65.18 51.69 56.92 61.22 71.40 68.77
13 68.41 82.39 97.95 85.81 91.11 100.00 99.15 95.96 72.52 89.90 54.70
14 84.22 51.71 91.95 63.05 47.57 83.58 78.83 98.36 83.72 96.70 95.67
15 71.70 46.45 91.31 39.98 40.51 99.20 87.32 81.48 76.42 69.49 76.77
16 38.17 84.34 87.95 89.96 78.31 100.00 99.60 100.00 58.84 78.78 67.54
OA (%) 57.54 41.93 71.11 50.02 36.02 69.58 66.53 71.12 72.75 85.08 85.54
AA (%) 52.66 58.43 79.56 63.73 53.36 81.28 80.94 80.99 72.58 78.30 76.68
Kappa (%) 51.81 35.65 67.77 44.14 30.33 65.98 62.71 67.40 69.05 83.04 83.58
  • 1
    The classes are as follows: 1. Alfalfa, 2. Corn-notill, 3. Corn-min, 4. Corn, 5. Grass-pasture, 6. Grass-tree, 7. Grass-pasture-mowed, 8. Hay-windrowed, 9. Oats,
    10. Soybeans-notill, 11. Soybeans-min, 12. Soybeans-clean, 13. Wheat, 14. Woods, 15. Bldg-grass-tree-drives, 16. Stone-stell-towers.
TABLE XXVII: Per-class classification accuracies of various methods on the Pavia University dataset.
Class1 MSDN [60] SSFCN [61] FullyContNet [59] SpectralFormer [20] HSIC-FM [62] SSGRN [63] CSIL [64] IDCN [65] CLOLN [66] SpatSIGMA HyperSIGMA
1 91.48 65.00 67.52 62.66 67.56 68.40 75.37 92.47 98.47 92.56 90.38
2 88.88 80.74 76.88 76.77 85.05 81.05 88.50 90.94 96.41 98.77 99.30
3 46.08 80.55 89.07 77.22 59.15 87.59 91.44 85.81 92.50 91.84 85.65
4 89.81 94.03 80.95 93.65 81.89 82.88 90.24 95.11 91.08 70.70 78.58
5 85.69 93.69 98.16 100.00 97.31 100.00 100.00 99.94 98.23 94.70 93.53
6 47.77 79.96 86.34 70.71 60.20 94.09 99.18 93.82 91.33 99.61 98.90
7 68.50 85.65 95.47 86.26 74.07 94.35 97.86 96.32 99.12 98.02 97.03
8 75.76 67.06 91.23 62.76 74.89 72.25 82.77 84.40 83.36 78.30 80.91
9 82.61 97.27 96.04 97.88 93.67 87.59 92.56 99.92 87.94 74.84 65.70
OA (%) 76.48 78.88 80.31 75.37 77.28 81.45 88.23 91.64 93.11 93.36 93.52
AA (%) 75.18 82.28 86.85 80.88 77.09 85.35 90.88 93.19 91.94 88.82 87.78
Kappa (%) 69.52 72.99 75.01 68.56 70.36 76.36 84.76 89.05 90.88 90.66 90.93
  • 1
    The classes are as follows: 1. Asphalt, 2. Meadows, 3. Gravel, 4. Trees, 5. Metal Sheets, 6. Bare Soil, 7. Bitumen, 8. Bricks, 9. Shadow.
TABLE XXVIII: Per-class classification accuracies of various methods on the HanChuan dataset.
Class1 MSDN [60] SSFCN [61] FullyContNet [59] SpectralFormer [20] HSIC-FM [62] SSGRN [63] CSIL [64] IDCN [65] CLOLN [66] SpatSIGMA HyperSIGMA
1 79.00 65.41 85.50 79.49 66.77 86.12 88.10 87.01 96.15 98.19 98.65
2 70.37 64.57 56.54 77.59 78.02 85.31 83.49 82.38 92.53 97.49 97.01
3 35.64 61.81 26.70 82.40 61.22 96.46 96.77 88.48 59.43 90.27 94.61
4 95.09 88.94 89.90 94.59 85.86 96.32 98.87 97.24 88.02 97.96 95.67
5 96.93 84.38 98.93 97.42 99.62 100.00 99.94 97.36 74.39 80.79 74.71
6 25.19 39.45 80.81 50.50 47.52 98.68 92.27 75.96 52.29 63.62 91.70
7 90.69 58.22 94.21 86.77 52.86 97.77 86.93 75.75 89.69 77.89 86.55
8 35.39 50.78 84.43 64.84 50.93 89.93 87.34 66.21 91.43 97.42 96.37
9 47.10 39.03 59.88 69.49 52.03 77.81 84.80 75.01 70.69 85.72 87.34
10 60.12 52.82 63.25 79.63 36.21 91.98 92.99 96.83 88.67 97.59 96.92
11 50.42 30.92 85.45 77.70 15.60 90.65 91.23 71.41 94.82 94.93 89.86
12 64.04 70.41 95.32 80.58 68.76 99.94 99.99 82.08 72.72 95.36 93.58
13 52.09 41.63 76.02 69.09 45.48 90.20 88.26 80.65 58.12 75.72 75.56
14 84.82 69.20 79.48 73.45 72.95 81.19 83.26 72.29 90.08 96.69 88.38
15 97.02 89.50 63.38 92.11 94.60 95.67 96.47 96.01 60.62 74.02 80.07
16 94.95 76.87 87.49 97.69 84.46 95.28 88.45 92.20 99.61 99.73 99.63
OA (%) 73.40 63.35 78.80 82.60 66.21 90.43 88.55 84.15 86.73 94.03 94.44
AA (%) 67.43 61.50 76.71 79.58 63.31 92.08 91.20 83.49 79.96 88.96 90.41
Kappa (%) 69.09 57.79 75.56 79.74 60.55 88.88 86.75 81.64 84.56 93.16 93.62
  • 1
    The classes are as follows: 1. Strawberry, 2. Cowpea, 3. Soybean, 4. Sorghum, 5. Water spinach, 6. Watermelon, 7. Greens, 8. Trees, 9. Grass, 10. Red roof,
    11. Gray roof, 12. Plastic, 13. Bare soil, 14. Road, 15. Bright object, 16. Water.
TABLE XXIX: Per-class classification accuracies of various methods on the HongHu dataset.
Class MSDN [60] SSFCN [61] FullyContNet [59] SpectralFormer [20] HSIC-FM [62] SSGRN [63] CSIL [64] IDCN [65] CLOLN [66] SpatSIGMA HyperSIGMA
1 73.82 80.74 78.29 94.51 66.32 96.73 93.05 94.27 95.20 98.16 98.10
2 89.69 78.79 60.18 91.06 85.91 79.86 95.90 92.06 80.97 76.91 61.49
3 64.26 67.70 85.21 79.22 75.81 91.26 93.96 81.48 98.00 95.27 97.22
4 90.11 81.41 73.83 93.12 73.81 85.84 91.47 94.74 99.48 99.71 99.71
5 66.95 60.53 92.27 89.18 68.98 99.91 98.87 92.02 68.49 75.39 74.17
6 88.81 76.67 68.22 89.66 82.36 69.17 93.98 94.35 96.32 96.67 98.30
7 34.58 48.47 26.87 64.63 44.84 62.22 79.68 73.56 91.40 94.90 94.75
8 34.49 34.17 81.17 62.40 58.40 98.52 98.43 73.12 54.66 89.38 90.82
9 87.52 76.45 56.46 95.01 74.54 89.46 94.27 94.75 89.31 98.78 95.30
10 73.75 58.24 66.86 65.30 79.34 58.49 93.03 79.07 84.71 97.34 96.76
11 41.96 39.76 51.95 57.98 38.83 68.51 85.52 79.78 66.22 88.16 87.71
12 63.45 60.67 67.14 68.95 59.34 88.85 94.19 77.52 69.82 77.51 89.24
13 68.56 50.10 56.42 65.17 54.05 74.81 89.42 66.16 80.22 81.72 92.23
14 64.67 70.85 70.70 79.28 73.51 84.84 96.10 93.22 78.09 94.88 85.45
15 90.44 78.99 93.03 95.55 92.40 100.00 99.96 96.34 44.52 69.35 59.97
16 94.65 81.44 66.56 90.75 87.12 98.98 97.51 93.95 97.24 95.94 96.50
17 89.00 84.21 92.71 90.44 90.02 99.96 99.27 95.17 88.21 96.73 82.55
18 90.21 65.99 82.70 90.05 67.97 97.12 98.58 93.79 77.60 87.03 83.30
19 65.36 52.88 66.13 87.60 60.16 77.47 91.04 90.70 70.47 83.77 84.64
20 88.86 63.27 93.02 94.83 74.19 99.33 99.63 93.74 67.00 80.40 84.95
21 59.57 66.82 99.53 91.94 86.25 100.00 100.00 94.32 66.89 83.65 79.08
22 86.60 79.98 93.03 93.65 68.44 99.42 98.20 95.25 56.61 76.60 85.84
OA (%) 78.55 71.62 67.12 85.33 70.47 82.19 91.86 89.19 87.89 94.35 94.87
AA (%) 73.06 66.51 73.81 83.20 71.03 87.31 94.64 88.15 78.25 88.10 87.18
Kappa (%) 73.45 65.67 61.34 81.76 64.73 78.21 89.89 86.59 84.83 93.04 93.69
  • 1
    The classes are as follows: 1. Red roof, 2. Road, 3. Bare soil, 4. Cotton, 5. Cotton firewood, 6. Rape, 7. Chinese cabbage, 8. Pakchoi, 9. Cabbage, 10. Tuber mustard,
    11. Brassica parachinensis, 12. Brassica chinensis, 13. Small Brassica chinensis, 14. Lactuca sativa, 15. Celtuce, 16. Film covered lettuce, 17. Romaine lettuce,
    18. Carrot, 19. White radish, 20. Garlic sprout, 21. Broad bean, 22. Tree.
TABLE XXX: Per-class classification accuracies of various methods on the Houston dataset.
Class MSDN [60] SSFCN [61] FullyContNet [59] SpectralFormer [20] HSIC-FM [62] SSGRN [63] CSIL [64] IDCN [65] CLOLN [66] SpatSIGMA HyperSIGMA
1 96.62 81.89 50.21 81.92 76.29 74.64 78.44 90.64 97.38 83.81 81.22
2 92.43 92.61 39.63 93.83 62.75 75.85 72.93 87.22 98.51 84.93 86.76
3 20.06 62.05 91.88 94.59 32.81 83.96 97.62 84.98 54.84 97.27 98.85
4 96.53 95.52 42.93 94.13 60.13 94.60 66.29 95.99 97.80 83.36 84.57
5 83.44 85.80 56.53 97.95 83.49 94.41 93.21 94.69 99.65 100.00 100.00
6 65.21 77.86 72.96 88.11 59.44 93.01 94.64 85.16 91.71 98.46 99.08
7 89.41 75.78 61.13 81.56 70.37 71.08 74.16 74.66 89.36 75.55 76.80
8 62.23 62.93 19.12 62.27 28.58 57.64 29.60 69.66 93.59 76.13 81.13
9 88.61 65.72 40.98 59.08 63.17 45.23 48.54 84.88 90.59 80.83 73.40
10 57.21 55.18 53.80 55.21 30.53 52.32 64.90 87.95 89.02 69.44 70.94
11 69.53 51.93 84.95 63.28 37.95 52.85 63.98 75.60 74.61 99.76 97.09
12 71.96 65.42 37.56 65.42 53.19 66.86 58.92 79.80 81.11 99.68 100.00
13 68.79 70.06 66.20 60.23 53.45 74.74 58.60 73.40 92.79 81.88 74.20
14 78.40 85.83 98.92 92.44 70.04 100.00 99.87 82.76 91.39 100.00 100.00
15 81.29 55.04 19.80 95.07 17.34 25.37 32.84 96.17 90.59 100.00 100.00
OA 72.18 72.39 51.07 77.21 54.43 68.62 66.11 85.34 85.95 87.33 86.80
AA 74.78 72.23 55.77 79.00 53.30 70.84 68.96 86.94 88.86 88.74 88.27
Kappa 69.93 70.12 47.00 75.32 50.80 65.99 63.21 84.11 84.78 86.30 85.72
  • 1
    The classes are as follows: 1. Healthy grass, 2. Stressed grass, 3. Synthetic grass, 4. Trees, 5. Soil, 6. Water, 7. Residential, 8. Commercial, 9. Road, 10. Highway,
    11. Railway, 12. Parking Lot 1, 13. Parking Lot 2, 14. Tennis Court, 15. Running Track.
TABLE XXXI: Per-class classification accuracies of various methods on the ZY1-02D dataset.
Class MSDN [60] SSFCN [61] FullyContNet [59] SpectralFormer [20] HSIC-FM [62] SSGRN [63] CSIL [64] IDCN [65] CLOLN [66] SpatSIGMA HyperSIGMA
1 85.39 72.02 54.78 41.49 90.45 77.45 79.58 93.59 87.05 88.49 90.57
2 81.28 70.24 68.69 32.99 68.25 77.89 89.49 91.47 87.35 92.01 94.86
3 99.82 94.04 94.28 96.34 80.16 83.80 97.83 96.43 97.46 98.62 99.24
4 94.60 86.27 77.00 50.59 89.91 78.52 89.79 95.07 49.06 98.10 92.25
5 93.83 89.24 90.32 89.22 88.67 77.09 97.75 94.92 94.24 97.30 98.20
[…]